OpenStudy (anonymous):

Solve (x-1)(x-2)=(k-1)(k-2), where k is a constant.

OpenStudy (anonymous):

Try expanding the L.H.S and R.H.S

OpenStudy (anonymous):

then...

OpenStudy (anonymous):

tell me !

OpenStudy (anonymous):

this is not the answer, the answer is x=k or x=3-k

OpenStudy (anonymous):

what is the solution, please help

OpenStudy (anonymous):

2x+2k-6=

OpenStudy (anonymous):

x should be =k that is clearly visible just compare the terms more of the brains should be applied on the second part

OpenStudy (anonymous):



OpenStudy (mathmale):

@vernonwan: as ╰☆╮Openstudier╰☆╮ has already suggested: Try expanding the L.H.S and R.H.S: (x-1)(x-2) = (k-1)(k-2) becomes what? Note that both expansions will have 2 as their 3rd term. Would you do this work now, please?

OpenStudy (anonymous):

\[x ^{2}-3x+2 = k ^{2} -3k +2\]

OpenStudy (mathmale):

Now cancel the 2 from both sides, and subtract (k^2-3k) from both sides next. You'll then have a quadratic equation in x. Try applying the quadratic formula to this quadratic equation.

OpenStudy (loser66):

As you said, x = 3-k is one of the solution I replace to test (3-k)^2 -3(3-k)+2 = 9 -6k+k^2-9 +6k+2 = k^2 +2 that is the left hand side . Question: do you really thing that k^2 +2 = k^2 +3k +2 ???? where k is a constant?

OpenStudy (loser66):

*think

OpenStudy (loser66):

I make that argument because your solution is not correct ( I mean the solution of x = 3-k). To me, the only one solution is x =k but you didn't accept it.

OpenStudy (anonymous):

thanks, I am trying , the answer is from my text book

OpenStudy (anonymous):

\[\frac{ x ^{2}-3x }{ k ^{2}-3k }=1\]

OpenStudy (mathmale):

Excuse my insistence, but would you please follow the suggestions I gave you earlier. You can solve this problem using the quadratic formula. Obtaining the solution comes first. Checking the solutions comes second (but is an important step).

OpenStudy (anonymous):

Your answer from the textbook is absolutely correct i worked it out

OpenStudy (anonymous):

then, \[x ^{2}-3x-k ^{2}+3k\] how to solve this

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

x^2 -3x = k^2 - 3k x^2 - k^2 = -3k + 3x (x-k)(x+k)= 3 (-k+x) now cancel out!!! x+k=3 x=k-3 yesssssssssss

OpenStudy (anonymous):

@mathmale i got it !

OpenStudy (mathmale):

\[x ^{2}-3x-k ^{2}+3k=0\]has the form of a quadratic equation, right? A reminder: k is a constant. You might want to re-write this as \[x ^{2}-3x+3k-k ^{2}=0\] and then (referring to ax^2 + bx + c = 0), identify your a as being equal to 1; b as being equal to -3; and c as being equal to 3k-k^2. Now substitute these a, b and c into the quadratic formula. If done correctly, this will give you the two solutions x=k and x=3-k.

OpenStudy (anonymous):

i had a simple method

OpenStudy (anonymous):

but your method is also good

OpenStudy (anonymous):

is my method correcct @mathmale

OpenStudy (anonymous):

@vernonwan did u get my method??

OpenStudy (mathmale):

x+k=3 x=k-3 is correct. What about the other solution, x=k? To be correct, any method must result in ALL solutions.

OpenStudy (anonymous):

i gave the method for the other sollution

OpenStudy (anonymous):

already

OpenStudy (anonymous):

yes thanks all of you